OpenStudy (margotlakus):

I need help with 3 more derivative problems if someone is willing to help!

OpenStudy (margotlakus):

\[y=\frac{ (x+1)(x-2) }{ (x-1)(x+2) }\] & \[x > 2\]

OpenStudy (margotlakus):

\[\sqrt{x}+\sqrt{y}=9\] y(1)=64 find y'(1) by implicit differentiation

OpenStudy (margotlakus):

\[\frac{ y }{ x-6y }=x^4-8\] at the point \[(1,\frac{ 7 }{ 41 })\]

OpenStudy (margotlakus):

\[\frac{ x^2 }{ 49 }+\frac{ y^2 }{ 16 }=1\] y(3)=3.61403161 y'(3)

OpenStudy (jackthegreatest):

do you know the quotient rule?

OpenStudy (margotlakus):

I do but I have to use logarithmic differentiation

OpenStudy (jackthegreatest):

ohh, i have not reached there in my math class yet : ( sorry

OpenStudy (tkhunny):

Okay, have you considered introducing the logarithm, then using a few log rules to manipulate?

OpenStudy (margotlakus):

I really have no idea what to do...

OpenStudy (tkhunny):

\(y = \dfrac{(x+1)(x-2)}{(x-1)(x+2)}\) Introduce the logarithm. \(log(y) = log\left[\dfrac{(x+1)(x-2)}{(x-1)(x+2)}\right]\) Use some log rules to expand or simplify and show your best implicit derivative methodology.

OpenStudy (margotlakus):

Is that the derivative or

OpenStudy (tkhunny):

Did you read it? I introduced the logarithm and told you to find the derivative. Does that mean it's the derivative or is there some work for you?

OpenStudy (margotlakus):

Oh, nevermind. Sorry my mind is very tired.

OpenStudy (tkhunny):

You can do it. Use your log rules. log(ab) = log(a)+log(b) and log(a/b) = log(a) - log(b). Those guys.

OpenStudy (margotlakus):

I need help with the other problems if anyone is willing

OpenStudy (mathmale):

Unsure of which problem you want help with. Avoid this ambiguity by posting just one problem at a time.